OpenStudy (anonymous):

evaluate the integral from 1 to 2 ((v^4+3v^5)/(v^3)) dv I am really struggling with this.

OpenStudy (anonymous):

your pic is so adorable

OpenStudy (anonymous):

I have gotten (v+3v^2)dv=[(1/2)v^3+v^4] I don't know what to do from there

OpenStudy (anonymous):

sweetie i am sorry i have an f on math!

OpenStudy (anonymous):

okay thank you though

zepdrix (zepdrix):

\[\Large\rm \int\limits_1^2 \frac{v^4+3v^5}{v^3}dv=\int\limits_1^2 \frac{v^4}{v^3}+\frac{3v^5}{v^3}dv\]\[\Large\rm \int\limits_1^2 v+3v^2\;dv=\frac{1}{2}v^2+\frac{3}{3}v^3|_1^2\]Hmm I think your powers might be off a little bit

OpenStudy (anonymous):

and then from there I just use the 1 and 2 in that?

zepdrix (zepdrix):

Yes :) Upper minus lower. If it helps to remember which comes first, think of it like old school subtraction: